# The Klein-Gordon equation on asymptotically Minkowski spacetimes: causal propagators

### v1

@@ 7. Radial point estimates over $C$
We proceed in two steps, first we show that we can estimate the right hand side of (7.6) by the right hand side claimed in Proposition 7.1 and then we construct $B_{0}$ from $B$ such that $B_{0}$ is elliptic at $\tau=m$.
Step 1:
We have to show that if $B_{1}u\in L^{2}$, then
$\lVert Bu\rVert\lesssim\lVert GP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,.$
By assumption $G$ is elliptic on ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q)$ and $G,x^{-1/2}Q$ are both of spatial order $\ell+1$. Therefore
$\lVert x^{-1/2}QP_{V}u\rVert\lesssim\lVert GP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,.$
By the previous lemma and a standard induction argument to replace $x^{-1}B$ by $B_{1}$, we obtain
$\lVert Bu\rVert\lesssim\lVert GP_{V}u\rVert+\lVert B_{1}u\rVert+\lVert u\rVert_{-N,-M}\,.$
Step 2:
Now we have to show that we can estimate
$\lVert B_{0}u\rVert\lesssim\lVert GP_{V}u\rVert+\lVert B_{1}u\rVert+\lVert u\rVert_{-N,-M}$
for some $B_{0}\in{}^{3\mathrm{sc}}\Psi^{s,\ell}$ with $m\in\operatorname{Ell}_{\operatorname{ff}}(B_{0})$ and ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{0})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G)$. If $B$ was elliptic at $m\in\overline{W^{\perp}}$, then we could just take $B_{0}=B$.
Let $Q^{\prime}=\operatorname{Op}_{L}(q^{\prime})\in{}^{\mathrm{sc}}\Psi^{0,0}$ with
$\hat{N}_{\operatorname{ff}}(Q^{\prime})(m)=2\sqrt{m(\ell+1/2)}\cdot\operatorname{Id}$
and ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q^{\prime})$ sufficiently small and set
$B^{\prime}\coloneqq B+x^{-\ell}Q^{\prime}(\operatorname{Id}-G_{\psi})\,.$
Then
$\hat{N}_{\operatorname{ff},\ell}(B^{\prime})$ $=2\sqrt{\tau(\ell+1/2)}\chi(\tau-m)\hat{N}_{\operatorname{ff}}(G_{\psi})+2\sqrt{m(\ell+1/2)}\hat{N}_{\operatorname{ff}}(\operatorname{Id}-G_{\psi})$
and
$\hat{N}_{\operatorname{ff},\ell}(B^{\prime})(m)=2\sqrt{m(\ell+1/2)}\cdot\operatorname{Id}\,.$
Therefore, $m\in\operatorname{Ell}_{\operatorname{ff}}(B^{\prime})$. By Corollary 5.13, we have that
$\lVert B^{\prime}u\rVert$ $\leq\lVert Bu\rVert+\lVert x^{-\ell}Q^{\prime}(\operatorname{Id}-G_{\psi})\rVert$
$\lesssim\lVert GP_{V}u\rVert+\lVert B_{1}u\rVert+\lVert u\rVert_{-N,-M}\,.$
@@
The argument is basically the same as in the above threshold case, but we can absorb $\lVert B_{1}u\rVert$ into $\lVert Eu\rVert$. More precisely, we can estimate
$\lVert B_{1}u\rVert\lesssim\lVert GP_{V}u\rVert+\lVert Eu\rVert+\lVert u\rVert_{-N,-M}$
by propagation of singularities. ∎
@@ 8. Construction of the causal propagators
    there exist $V_{\pm}\in S^{-1}(\mathbb{R}^{n}_{z})$ and $C>0$ such that for all $t>C$ and $\lvert z/t\rvert<C$,
    $V(t,z)-V_{+}(z)\in{}^{3\mathrm{sc}}\Psi^{1,-1}\,,$
    and for all $t<-C$ and $\lvert z/t\rvert<C$,
    $V(t,z)-V_{-}(z)\in{}^{3\mathrm{sc}}\Psi^{1,-1}\,.$
@@
In the case that the Hamiltonians $H_{V_{\pm}}$ have discrete spectrum we have to strengthen the assumptions on the decay of $V-V_{\pm}$: To simplify our exposition, we make the further assumption that, as $t\to\pm\infty$, for $V_{\pm}$ the asymptotic values of $V$ in Section 3.4, we have an additional order of spacetime decay,
@@
Let $V\in\rho_{\operatorname{mf}}{}^{3\mathrm{sc}}\Psi^{1,0}$ satisfying (1), (2), and (3). Assume that as $t\to\pm\infty$ we have
(8.1) $V-V_{\pm}(z)\in{}^{3\mathrm{sc}}\Psi^{1,-2}(\mathbb{R}^{n+1}).$
Then the mapping
@@ 8.1. Assuming no bound states
We start by proving Theorem 8.2. The assumption that there are no eigenvalues in $(-\infty,m^{2})$ implies that
@@
Indeed, recalling $(\widehat{P_{V}})_{\operatorname{ff}}(\tau)=\tau^{2}-(\Delta_{z}+m^{2}+V_{+}(z))$, we see that for $|\tau|<m$, $(\widehat{P_{V}})_{\operatorname{ff}}(\tau)$ is scattering elliptic. Thus (see Section 4.6), for $|\tau|<m$, we have that $\tau\in\operatorname{Ell}_{\operatorname{ff}}(P_{V})$ if and only if $\tau^{2}-(\Delta_{z}+m^{2}+V_{+})$ is in fact invertible. Thus, ellipticity of $P_{V}$ at $\tau\in W^{\perp}\cap(-m,m)$ is equivalent to the non-existence of a bound state of $\Delta+V_{+}$ with energy $E=\tau^{2}-m^{2}$. Thus in this section we will have (8.3) at both $\mathrm{NP}$ and $\mathrm{SP}$.
For $u\in\mathcal{S}(\mathbb{R}^{n+1})$, define
[…]
where $c(t)$ is the minimum of $\sigma(\Delta+m^{2}+V(t,z))$. If there are no bound states of $V_{-}$, then $\sigma(\Delta+m^{2}+V_{-})\geq c_{0}>0$, so by the Kato–Rellich theorem, there is $t_{0}$ such that for any $t<t_{0}$, $E_{u}(t)\geq(c-\delta)\|u\|^{2}$, which is to say in particular that $E_{u}(t)\geq 0$ for $t$ sufficiently small. The same goes for $t\to+\infty$, so $E_{u}(t)\geq 0$ for any $|t|>t_{0}^{\prime}>0$.
We first prove that $P_{V}:\mathcal{X}^{s,\boldsymbol{\ell}_{+}}\to\mathcal{Y}^{s-1,\boldsymbol{\ell}_{+}+1}$ is a Fredholm mapping. Again this reduces to showing the analogue of (2.33), namely that for any $N,M,s\in\mathbb{R}$, there is $C>0$ such that, provided all quantities are finite,
@@
$\lVert u\rVert_{1-s,-1-\boldsymbol{\ell}_{+}}$ $\leq C\left(\lVert P_{V}u\rVert_{-s,-\boldsymbol{\ell}_{+}}+\lVert u\rVert_{-N,-M}\right),$
Again, these estimates imply the Fredholm property by a standard argument.
To obtain them we again argue as in Section 2.8. We choose an open cover $O_{1},O_{2},O_{3},O_{4}$ of the compressed cotangent bundle ${}^{\mathrm{sc}}\dot{\overline{T}}^{*}X$. Note here that, due to our assumption of no bound states, ${}^{3\mathrm{sc}}\operatorname{Char}P_{V}={}^{3\mathrm{sc}}\operatorname{Char}P_{0}$.
@@
    ±$\mathcal{R}_{\pm}^{p}\subset O_{1}\subset\{\boldsymbol{\ell}_{+}=-1/2+\epsilon\}$, in particular ±$\pm m\in O_{1}\cap W^{\perp}$ over $\mathrm{SP}$.
@@
    ±$\mathcal{R}_{\pm}^{f}\subset O_{2}\subset\{\boldsymbol{\ell}_{+}=-1/2-\epsilon\}$, in particular ±$\pm m\in O_{2}\cap W^{\perp}$ over $\mathrm{NP}$.
@@
    $O_{2}\setminus\mathcal{R}^{f}$ is controlled along ${}^{\mathrm{sc}}\!H$ by $O_{3}$, and
@@
In particular, $O_{1}$ and $O_{2}$ will be small neighborhoods around $\mathcal{R}$ over $\mathrm{SP}$ and $\mathrm{NP}$ respectively.
@@
Keeping this in mind, we now choose a collection $B_{1},B_{2},B_{3},B_{4}\in{}^{3\mathrm{sc}}\Psi^{0,0}$ (in fact in ${}^{\mathrm{sc}}\Psi^{0,0}$) with
(8.11) ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{i})\subset O_{i}$
as in Section 2.8. Due to the overlap just described, the conditions ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{i})\subset O_{i}$ are ambiguous, in that they do not indicate the behavior of the $B_{i}$ at the fiber equator above $C$. To clarify this, we take the approach that we require that
@@
meaning each point in $\dot{C}_{3\mathrm{sc}}[X;C]X$ is is in the elliptic set of some $B_{i}$ (4.30). Over $\mathrm{SP}$ or $\mathrm{NP}$, since ellipticity at $\tau_{0}\in W^{\perp}$ requires ellipticity of the symbol out to fiber infinity over $\tau=\tau_{0}$, the condition is incompatible with the $B_{i}$ forming a partition of unity. Note however that any collection $B_{i}$ satisfying (8.12) gives the “same control” as a partition of unity, namely, for any $N,M\in\mathbb{R}$, and any $s,\ell$,
@@
To construct such $B_{i}$ which satisfy both (8.11) and (8.12) we can use cutoff functions and the expressions for the $O_{i}$ above. For example, we can take $B_{1}=\operatorname{Op}_{L}(\chi_{1})$ with $\chi_{1}\colon{}^{\mathrm{sc}}\overline{T}^{*}X\longrightarrow\mathbb{R}$ identically $1$ on sets of the forms (8.9) and (8.10) with the $\epsilon$ replace by a smaller $\epsilon^{\prime}>0$, and with support in the union of (8.11). Such $B_{1}$ has $\hat{N}_{\operatorname{ff}}(B_{1})(\tau)=\operatorname{Id}$ for $\tau\in[-m-\epsilon^{\prime},-m+\epsilon^{\prime}]\cup[m-\epsilon^{\prime},m+\epsilon^{\prime}]$ and indeed $\mathcal{R}^{p}\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(B_{1})$. We define $B_{2}$ and $B_{3}$ similarly.
Deducing the estimates in (8.5) for $u$ using the $B_{i}$ now follows exactly along the lines of the deduction in Section 2.8, by obtaining estimates exactly as in (2.34) - (2.37) with $P_{V}$ replacing $P_{0}$. Thus, again, $u\in\mathcal{X}^{s,\boldsymbol{\ell}_{+}}$, then $u$ is above threshold near ±$\mathcal{R}_{\pm}^{p}$, now in the $3\mathrm{sc}$-sense. That is:
  1. (1)
    Putting together Proposition 2.11 and Proposition 7.1, we obtain (2.34) (with $P_{V}$ replacing $P_{0}$) with the same $\boldsymbol{\ell}^{\prime}<\boldsymbol{\ell}_{+}$ and $-1/2<\boldsymbol{\ell}^{\prime}<\boldsymbol{\ell}_{+}$.
  2. (2)
    Again, since ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{3})$ is controlled by ${}^{3\mathrm{sc}}\!\operatorname{Ell}(B_{1})$, by Proposition 6.1 we have (2.35) (with $P_{V}$).
  3. (3)
    By Proposition 2.13 and Proposition 7.2 we have the below threshold estimate for $B_{2}$ in (2.36), and
  4. (4)
    We have the $3\mathrm{sc}$-elliptic estimates so by Proposition 4.20 we obtain (2.37) (with $P_{V}$).
Since we control $u$ by the $B_{i}u$ via (8.13) we again obtain the Fredholm estimates in (8.5), the second estimate being deduced by applying the above and below threshold estimates to the opposite radial sets. We again can absorb the $\boldsymbol{\ell}^{\prime}$ term on the left by using (2.38). Thus the operator in (8.2) is Fredholm.
@@
must, thanks to the above threshold radial points estimates, in fact be rapidly decaying along with all their derivatives to the past. Using the energy $E_{u}(t)$ and a standard Grönwall argument gives that $u(t,z)\equiv 0$ for $t\gg 0$. Indeed, $dE_{u}(t)/dt\leq k(t)E_{u}(t)$ where $k(t)=O(1/|t|)^{2}$ as $t\to-\infty$. Since $E_{u}(t)\to 0$ as $t\to-\infty$ the differential Grönwall inequality shows that $E_{u}(t)\equiv 0$ for $t\ll 0$. By the positivity of the energy, we obtain that $u(t)=0$ for $t\ll 0$. Then uniqueness of solutions to the Cauchy problem shows $u\equiv 0$ globally.

### v2

@@ 7. Radial point estimates over $C$
The proof follows from exactly the same arguments as the propagation of singularities result, Proposition 6.2 follows from the explicit construction in Lemma 6.12. ∎
@@
Again the argument is very similar, but we can decrease the decay order $\ell-k/2$ in $\lVert B_{1}u\rVert_{s-k/2,\ell-k/2}$ only to $\ell^{\prime}$ for $\ell^{\prime}>-1/2$ and then we use Proposition 5.14 to estimate
$\lVert B_{1}u\rVert_{s^{\prime},\ell^{\prime}}\lesssim\lVert B_{1}G_{\phi_{1}}u\rVert_{s^{\prime},\ell^{\prime}}+\lVert G^{\prime}P_{V}u\rVert_{s-2,\ell}+\lVert B^{\prime}u\rVert_{s-r_{1}-1,\ell-r_{2}-1}+\lVert u\rVert_{-N,-M}\,.$
@@ 8. Construction of the causal propagators
    $V$ is asymptotically static of order $1$ in the sense of Definition 5.13.
@@
In the case that the Hamiltonians $H_{V_{\pm}}$ have discrete spectrum we have to strengthen the assumptions on the decay of $V-V_{\pm}$, namely we assume to have an additional order of spacetime decay.
@@
Let $V\in\rho_{\operatorname{mf}}{}^{3\mathrm{sc}}\Psi^{1,0}$ be asymptotically static of order $2$ and satisfies (2) and (3). Then the mapping
@@ 8.1. Assuming no bound states
In this section, we prove Theorem 8.2. The assumption that there are no eigenvalues in $(-\infty,m^{2})$ implies that
@@
Indeed, the assumption implies directly that $\hat{N}_{\operatorname{ff}_{\pm}}(P_{V})(\tau)=\tau^{2}-(\Delta_{z}+m^{2}+V_{\pm}(z))$ is invertible for $\lvert\tau\rvert<m$, which implies the ellipticity statement. (See Section 4.6.) Thus in this section we will have (8.2) at both $\mathrm{NP}$ and $\mathrm{SP}$.
We first prove that $P_{V}:\mathcal{X}^{s,\boldsymbol{\ell}_{+}}\to\mathcal{Y}^{s-1,\boldsymbol{\ell}_{+}+1}$ is a Fredholm mapping. This reduces to showing the analogue of (2.33), namely that for any $N,M,s\in\mathbb{R}$, there is $C>0$ such that, provided all quantities are finite,
@@
$\lVert u\rVert_{1-s,-1-\boldsymbol{\ell}_{+}}$ $\leq C\left(\lVert P_{V}u\rVert_{-s,-\boldsymbol{\ell}_{+}}+\lVert u\rVert_{-N,-M}\right).$
To obtain these estimates we again argue as in Section 2.8. We choose an open cover $O_{1},O_{2},O_{3},O_{4}$ of the compressed cotangent bundle ${}^{\mathrm{sc}}\dot{\overline{T}}^{*}X$. Note here that, due to our assumption of no bound states, ${}^{3\mathrm{sc}}\operatorname{Char}P_{V}={}^{3\mathrm{sc}}\operatorname{Char}P_{0}$. With $c_{1},c_{2}>0$,
@@
    ±$\mathcal{R}_{\pm}^{p}\subset O_{1}\subset\{\boldsymbol{\ell}_{+}=-1/2+c_{1}\}$, in particular ±$\pm m\in O_{1}\cap W^{\perp}$ over $\mathrm{SP}$.
@@
    ±$\mathcal{R}_{\pm}^{f}\subset O_{2}\subset\{\boldsymbol{\ell}_{+}=-1/2-c_{2}\}$, in particular ±$\pm m\in O_{2}\cap W^{\perp}$ over $\mathrm{NP}$.
@@
    $O_{2}\setminus\mathcal{R}^{f}$ is controlled along ${}^{\mathrm{sc}}\!H_{p}$ by $O_{3}$, and
@@
In particular, $O_{1}$ and $O_{2}$ are neighborhoods of $\mathcal{R}^{p}$ and $\mathcal{R}^{f}$ respecitively, but over $\mathrm{NP}$ and $\mathrm{SP}$ we restrict their size for convenience.
@@
Keeping this in mind, we now choose a collection $B_{1},B_{2},B_{3},B_{4}\in{}^{3\mathrm{sc}}\Psi^{0,0}$ (in fact in ${}^{\mathrm{sc}}\Psi^{0,0}$) analogous to the $B_{i}$ in the scattering construction in Section 2.8. However, recalling that
$\operatorname{Ell}_{\operatorname{ff}}(B)\neq\varnothing\implies\overline{W^{\perp}}=\operatorname{WF}^{\prime}_{\operatorname{ff}}(B),$
we cannot have the ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}$ subordination condition (2.34) on the nose. In fact, this condition is only a convenient means of relating the elliptics sets of the $B_{i}$ to the assumed Hamiltonian dynamics. Here we instead take the approach that the $B_{i}$ will have ellip
Due to the overlap just described, the conditions ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{i})\subset O_{i}$ are ambiguous, in that they do not indicate the behavior of the $B_{i}$ at the fiber equator above $C$. Here we instead require that
@@
meaning each point in $\dot{C}_{3\mathrm{sc}}[X;C]$ is in the elliptic set of some $B_{i}$ (4.30). Any collection $B_{i}$ satisfying (8.9) gives the “same control” as a partition of unity, namely, for any $N,M\in\mathbb{R}$, and any $s,\ell$,
@@
As a replacement for the condition ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{i})\subset O_{i}$, we assume that, for some $\delta>0$
${}^{3\mathrm{sc}}\!\operatorname{Ell}(B_{i})$ $\subset O_{i},$
$\operatorname{WF}_{\operatorname{ff}}^{\prime}(B_{i}G_{\psi})$ $\subset O_{i}\mbox{ for }\psi\in C_{c}^{\infty}(\mathbb{R})\mbox{with }\operatorname{supp}\psi\subset(-\delta,\delta),\mbox{ and }$
$\operatorname{WF}_{\bullet}^{\prime}(B_{i})\cap{}^{3\mathrm{sc}}\operatorname{Char}(P_{V})$ $\subset O_{i}\mbox{ for }\bullet\in\{\operatorname{fib},\operatorname{mf}\}.$
To construct such $B_{i}$ which satisfy both these conditions and (8.9) we can use cutoff functions and the expressions for the $O_{i}$ above. For example, we can take $B_{1}=\operatorname{Op}_{L}(\chi_{1})$ with $\chi_{1}\colon{}^{\mathrm{sc}}\overline{T}^{*}X\longrightarrow\mathbb{R}$ identically $1$ on sets of the forms (8.7) and (8.8) with the $\epsilon$ replace by a smaller $\epsilon^{\prime}>0$, and with support in the union of (2.34). Such $B_{1}$ has $\hat{N}_{\operatorname{ff}}(B_{1})(\tau)=\operatorname{Id}$ for $\tau\in[-m-\epsilon^{\prime},-m+\epsilon^{\prime}]\cup[m-\epsilon^{\prime},m+\epsilon^{\prime}]$ and indeed $\mathcal{R}^{p}\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(B_{1})$. We define $B_{2}$ and $B_{3}$ similarly.
We now follow the process in Section 2.8, meaning we begin with above threshold estimates to bound $B_{1}u$, propagation estimates to bound $B_{3}u$ (in terms of $B_{1}u$), below threshold estimates to bound $B_{2}u$ (in terms of $B_{3}u$) and then elliptic estimates for $B_{4}u$. For example, for $B_{1}u$, Proposition 7.1 implies that, for any $\ell^{\prime}$ with $-1/2<\ell^{\prime}<\ell$,
(8.11) $\lVert B_{1}u\rVert_{s,\ell}\leq C\left(\lVert\tilde{B}u\rVert_{s-1,\ell^{\prime}}+\lVert P_{V}u\rVert_{s-1,\ell+1}+\lVert u\rVert_{-N,-M}\right)\,.$
where $B\in{}^{3\mathrm{sc}}\Psi^{0,0}$, ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{1}G_{\psi})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(\tilde{B})$. To see that this follows from the proposition, note first that all the terms in the above threshold estimate involving $P_{V}u$ can be bounded by the global norm of $P_{V}u$. Similarly, the terms on the right hand side involving $G_{\phi}u$ and $B^{\prime}u$ are both bounded by the $\tilde{B}u$ term in (8.11). Similarly,
(8.12) $\lVert B_{3}u\rVert_{s,\ell}\leq C\left(\lVert B_{1}u\rVert_{s,\ell}+\lVert P_{V}u\rVert_{s-1,\ell+1}+\lVert\tilde{B}^{\prime}u\rVert_{s-1,\ell-1}+\lVert u\rVert_{-N,-M}\right)\,,$
where $\tilde{B}^{\prime}$ comes from the $B^{\prime}$ term in Proposition 6.2. For the below threshold estimate we have
(8.13) $\lVert B_{2}u\rVert_{s,\ell}$ $\leq C\big{(}\lVert B_{3}u\rVert_{s,\ell}+\lVert P_{V}u\rVert_{s-1,\ell+1}+\lVert u\rVert_{-N,-M}\big{)}\,.$
Note here that the term from Proposition 7.2 with $B^{\prime}u\in H_{\mathrm{sc}}^{s-r_{1}-1,\ell-r_{2}-1}$ does not appear as it can be iterated and absorbed into the residual term.
Putting these estimates together with the elliptic estimates (Proposition 4.20) and using (2.39) to remove the $s-1,\ell-1$ term yields the global estimate which in turn gives the Fredholm result.
@@
must, thanks to the above threshold radial points estimates, be rapidly decaying along with all their derivatives to the past. This allows us to use a standard argument to show that the energy must be zero.
Indeed, we use the energy functionsl $E_{u}(t)$ with
[…]
where $c(t)$ is the minimum of $\sigma(\Delta+m^{2}+V(t,z))$. If there are no bound states of $V_{-}$, then $\sigma(\Delta+m^{2}+V_{-})\geq c_{0}>0$, so by the Kato–Rellich theorem, there is $t_{0}$ such that for any $t<t_{0}$, $E_{u}(t)\geq(c-\delta)\|u\|^{2}$. By a standard Grönwall argument, this gives that $u(t,z)\equiv 0$ for $t\gg 0$. Indeed, $dE_{u}(t)/dt\leq k(t)E_{u}(t)$ where $k(t)=O(1/|t|)^{2}$ as $t\to-\infty$. Since $E_{u}(t)\to 0$ as $t\to-\infty$ the differential Grönwall inequality shows that $E_{u}(t)\equiv 0$ for $t\ll 0$. By the positivity of the energy, we obtain that $u(t)=0$ for $t\ll 0$. Then uniqueness of solutions to the Cauchy problem shows $u\equiv 0$ globally.